OpenStudy (ray1998):

Please help! Fan and medal will be given. Write an explicit formula for the following sequence. Find the fifteenth term. 5, 6, 7, 8, 9,.... What is an explicit formula for the given sequence? an = Thank you in advance!

OpenStudy (amilapsn):

\(a_1=5\\ a_2=6\\ a_3=7 \) and so on... then, \( a_n=?\)

OpenStudy (yttrium):

First you need to find they're common difference (we will denote them as d) Example: 6-5 = 1 7-6 = 1 8-7 = 1 9-8 = 1 so from that, we can observe that d = 1 Second, we will apply the formula for sequence\[a_{n} = a _{1} + d(n-1)\] By substitution, we will get\[a_{n} = 5 + 1 (n-1)\] or simply \[a_{n} = n+4\]

OpenStudy (yttrium):

oh no, it should be *their not they're

OpenStudy (mrnood):

glad to see that the correct use of grammar is not dead!